OpenStudy (anonymous):

if points A,B,C AND Dare arranged on a line in that order if AC=13 BD=14 and AD=21then what is BC

hartnn (hartnn):

hey @destiny123 ! \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (jhannybean):

\[\large \cfrac{a}{b} = \cfrac{0.625}{1}\]\[\large \cfrac{b}{a} = \cfrac{1}{0.625}\]

hero (hero):

\[\frac{a}{b} = \frac{625}{1000}\] \[\frac{b}{a} = \frac{1000}{625}\] Simplify that

OpenStudy (jdoe0001):

hmm, "somebody" edited the original question

hero (hero):

@destiny123 did

hero (hero):

@destiny123, you shouldn't edit your original question. You make all our responses look foolish.

hero (hero):

@shadowfiend, this is why "Edit Your Question" feature should be limited to three or four minutes.